# Defend Data Poisoning Attacks on Voice Authentication

### v1

@@ 4.1 Design Philosophy
In practice, it is actually quite challenging to calculate similarities between the raw audio files due to the differences in the audio length and content. In order to obtain meaningful analysis results, most voice recognition systems preprocess audio files by removing mute portions and converting them to 64-dimensional Fbank coefficients. In our experiments, each raw audio file is represented as a $160\times 64$ array. Then, we conduct the t-distributed stochastic neighbor embedding (t-SNE) on the normalized audio files for both normal accounts and attacked accounts as follows. Specifically, both normal accounts and attacked accounts need to provide $N$ audio files to the voice recognition system for training. For a normal account, we randomly split the $N$ audio files into two groups, each with $\frac{N}{2}$ files. Then, we perform the t-SNE analysis on the two groups to identify their similarity. As for the attacked account, we group $\frac{N}{2}$ victim’s audio files together and compare them with $\frac{N}{2}$ attacker’s files using the t-SNE. The comparison results of the two cases are shown in Figure 4 (a) and (b), respectively. As we can see, there is not an obvious difference in the data distribution between the normal account and the attacked account. That means comparing raw audio files may not be an effective way to detect the attackers.
@@ 5 Experimental Studies
In the experiments, we use two datasets: LibriSpeech [46] and VCTK [47]. The LibriSpeech and VCTK datasets contain audio files from 2484 and 108 people, respectively. For each person, we randomly select 10 audio files, and most of the audio files are several seconds long. 50% of LibriSpeech dataset users are used to train Deep Speaker models. 10 Deep Speaker models are trained by varying the weight initialization and the set of users being attacked. We evaluate the scenarios when the percentages of attacked users are 5% or 10%. For the Guardian network, 10 interleaved feature vectors are generated for each account and the value $K$ is set to 11 for the KNN classifier. To choose the optimal $K$ value, we first plot the curve of the error rate and $K$ with $K$ varying from 1 to 30. Then, according to the graph, we select the $K$ value with the projected minimum error rate.
@@
All the experiments were conducted on a computer with Intel i9-10900X CPU@3.7GHz, NVIDIA GeForce RTX 3090 GPU, and 64GBs of memory. In all the experiments, our Guardian model takes about 25 minutes for training, and only 3.3 seconds for validating a user. In what follows, we focus on evaluating its effectiveness in terms of prediction accuracy and recall as defined in the following equations.
@@
5.1 Varying the Percentage of Attackers
In the first round of experiments, we aim to compare the performance of our Guardian network with conventional machine learning approaches and the latest defense mechanism that used only fully connected layers (denoted as ”FC” in the figures) [16]. Table I shows the prediction accuracy and recall when there are 5% and 10% of poisoned user accounts. In both cases, we observe that our proposed Guardian network has achieved around 95% detection accuracy and recall, whereas other approaches have less than 60% accuracy and lower recall. The recall of the Guardian is also much higher than the other three approaches. These can be attributed to the series of techniques adopted by the Guardian network. The results clearly demonstrate the significant benefits of bias reduction, input augmentation, convolutional layers and ensemble learning.
TABLE I: Datasets with 5% and 10% Attacker Accounts
Poisoning Ratio Method Accuracy Recall
5% Attacker Accounts SVM 0.543 0.754
KNN 0.520 0.583
FC 0.639 0.804
Guardian 0.950 0.953
10% Attacker Accounts SVM 0.488 0.452
KNN 0.520 0.528
FC 0.529 0.588
Guardian 0.939 0.910
@@ 5.2 Effect of Bias Reduction
In this round of experiments, we are interested in examining the effect of our bias reduction technique alone. Table II shows the detection accuracy and recall with and without the bias reduction technique. The percentage of poisoned accounts is set to 5%. The version without bias collects training data from only one Deep Speaker model and has the ratio of the attacked accounts to the normal accounts as 1:20. The version with bias reduction used 4 Deep Speaker models and increased the ratio of the attacked accounts to the normal accounts to 1: 5.
TABLE II: Effect of Bias Reduction
Bias Reduction Accuracy Recall
No 0.986 0.266
Yes 0.943 0.952
@@ 5.3 Effect of Input Augmentation
We also evaluate the effectiveness of our proposed input augmentation, which interleaves a pair of feature vectors from the same account. Specifically, we modify the Guardian network to take the 512-dimensional feature vectors generated by the Deep Speaker model as input. We still keep the bias reduction and ensemble learning techniques for the 512-D Guardian network so as to single out the effect of the input augmentation. We compare this modified Guardian network with the Guardian network that has the input augmentation and takes 1024-dimensional feature vectors.
Table III reports the performance of these two networks. We can observe that a significant performance improvement has been achieved by the input augmentation technique for both the accuracy and recall. Specifically, the accuracy has been increased from 66% to 95%, and the recall has been increased from 35% to 95%. This is because when 512-dimensional feature vectors are used, the convolutional layers only study the spatial relationship in a single vector. The interleaved 1024-dimensional feature vectors give the convolutional layers an opportunity to compare the features from the attacker and victim dimension by dimension, and hence lead to better classification capabilities.
TABLE III: Effect of Input Augmentation
Input Type Accuracy Recall
512-D (Original) 0.663 0.350
1024-D (Augmented) 0.944 0.952
@@ 5.4 Effect of Ensemble Learning
In Table IV, we show the performance of two versions of the Guardian network with and without ensemble learning. The one without ensemble learning conducts only one prediction by taking a single 1024-dimensional feature vector from a randomly selected pair of 512-dimensional feature vectors from a new user account. The one with the ensemble learning conducts 10 predictions from 10 randomly selected pairs of the 512-dimensional feature vectors from the same user account.
@@
TABLE IV: Effect of Ensemble Learning
Ensemble Learning Accuracy Recall
No 0.866 0.458
Yes 0.944 0.952
@@ 5.5 Effect of Attackers’ Voices
In the previous experiments, both the victims and attackers are chosen randomly, which means the voices of the victim and the attacker could be relatively similar (e.g., of the same gender) or very different. The attack has been successful in both cases. We are interested in finding out if an attacker who has a similar voice as the victim would impose more challenges on Guardian. In this experiment, we use the Deep Speaker and Guardian models trained for Section 5.3. Then, we test Guardian against two kinds of attackers. Specifically, we select attackers of the same gender as the victims to simulate the similar voice scenario. Table V compares the detection accuracy of Guardian under different attack scenarios. We can observe that the Guardian’s performance is not affected by the similarity between attacker and victim’s voices.
TABLE V: Effect of Attacker’s Voices
Attacker’s Voices Accuracy Recall
Mixed 0.944 0.952
Same Gender 0.942 0.967
Different Gender 0.942 0.976
5.6 Model Transferability
In the last round of experiments, we are interested in learning the transferability of the Guardian model. We trained the Guardian model using the LibriSpeech dataset, and then we tested its effectiveness using the data from the LibriSpeech and VCTK dataset, respectively. Table VI shows the performance comparison. It is exciting to see that our Guardian model transfers well to a new data distribution. Observe that the detection accuracy in both datasets is similarly high, i.e., 95%. The recall in the new dataset is slightly lower but still around 80%. This is expected since the data distribution in VCTK is different from LibriSpeech that is used to train the Guardian model. The reason that the Guardian model has relatively good transferability is likely a combined effect of the multiple learning techniques adopted by the Guardian, especially the input augmentation which may play a major role in the transferability.
TABLE VI: Model Transferability
Dataset Accuracy Recall
Trained on LibriSpeech 0.944 0.952
Tested on VCTK 0.961 0.800
@@ 6 Security Analysis
We now analyze potential attacks to our Guardian model. The first scenario that may seem to challenge the effectiveness of the Guardian model is if the attacker has a similar voice to the victim. The similarity here simply refers to how the voice sounds to the human ears. Fortunately, the voice recognition models have much better voice identification capabilities than humans. They already have very high recognition accuracy among a large number of users whereby users of similar voices inevitably exist. In our experiments (Section 5.5), we also simulated such a scenario by selecting same-gender attacker and victims since their voices would be more similar than those from different genders. The experimental results show that our Guardian model performs similarly well in both cases. That means attackers with similar voices as the victim do not possess any advantages.
TABLE VII: Voice Recognition Accuracy Under Fake Voice Attacks
@@
Another scenario is when the attacker knows the existence and architecture of our Guardian model and attempts to take advantage of that. The attackers may try to use some existing Text-to-Speech (TTS) techniques to generate voice files that mimic the victim’s voices so as to fool the voice authentication model and the Guardian model. However, even if the fake voices are successfully generated to fool the voice authentication model and the Guardian model, they will not be able to escape from the fake voice detector which is good at distinguishing fake voices from authentic human voices. We have conducted the following experiments to validate our conjecture. First, we used a well-known repository called Real-Time Voice Cloning [49] to generate fake voices. This repository is an implementation of SV2TTS [50]. There are three different neural network structures in this system, and we use the pre-trained models they provided during our experiments. We chose 1166 users in LibriSpeech [46] dataset. Keeping the same settings as that in the previous experiments, each user has 10 original voice files, and 5% of user accounts (i.e., victims) contain fake voice files. Each victim account has 5 original voice files and 5 fake voice files generated by the Real-Time Voice Cloning system. Table VII shows the accuracy of voice recognition before and after the data poisoning attack. Before the attack, the overall voice recognition accuracy is around 0.991. After some user accounts being injected with fake voices, the recognition accuracy of both normal accounts and victim accounts still stay at a very high level as shown in the table. That means the fake voices have successfully fooled the voice recognition model. Next, we check how the Guardian model reacts to the fake voices. Table VIII shows the detection accuracy and recall. From the low recall rate, we can see that the Guardian model misses the majority of fake voices.
TABLE VIII: Guardian with 5% Fake Voice Users
Poisoning Ratio Accuracy Recall
5% Fake Voice 0.967 0.124
@@
The results from Table VII and VIII indicate that both the voice recognition model and the Guardian model are unaware of the fake voice attack. However, there is an easy way for the service providers to filter out such fake voices by employing the existing fake voice detectors such as Deep Sonar [48]. This detector is based on monitoring neuron behaviors of the voice recognition system to discern synthesized fake voices. In order to observe the performance of Deep Sonar in our system, we trained a brand new Deep Sonar network for the Deep Speaker using the LibriSpeech dataset. Table IX shows the high detection accuracy and recall achieved by Deep Sonar with respect to the fake voice attack. This means Deep Sonar is a very effective tool to detect fake voices and is powerful enough to defend against the fake voice attack.
@@
Note that fake voice detectors are a good complement to the overall voice authentication system, but cannot replace the function of the Guardian model since fake voice detectors are only versed at identifying manipulated voices. The second row in Table IX shows nearly zero recall rate with respect to poisoning attacks using human voices. In other words, Deep Sonar is not able to filter out authentic human voices when the attacker directly injected their real voices in the victim’s account like the targeted data poisoning attack in our case.
TABLE IX: DeepSonar under Different Types of Attacks
Attack Name Accuracy Recall
5% Fake Voice Attack 0.941 0.939
5% Poisoning Attack 0.891 0.05
@@
Finally, we discuss the scenario when attackers attempt to apply the idea of Generative Adversarial Networks (GAN) [51] to produce voice files that can fool the voice authentication model, the fake voice detector, and the Guardian model. However, it would be extremely challenging to implement such an attack. Recall that both the fake voice detector and the Guardian model hide behind the voice recognition model. As the attacker does not have the specific parameters of any of these three models, they may train their own system with the same structures. Since all the three models are deep neural networks, the back propagation process is very complicated. In our trial with known system parameters, we are still not able to make such training converge. Even if the attackers managed to complete the local training, it is unclear if the locally generated fake voices will be sufficiently effective to fool the real models used by their targeted service providers.

### v2

@@ 4.1 Design Philosophy
In practice, it is actually quite challenging to calculate similarities between the raw audio files due to the differences in the audio length and content. In order to obtain meaningful analysis results, most voice recognition systems preprocess audio files by removing mute portions and converting them to 64-dimensional Fbank coefficients. In our experiments, each raw audio file is represented as a $160\times 64$ array. Then, we conduct the t-distributed stochastic neighbor feature vector (t-SNE) on the normalized audio files for both normal accounts and attacked accounts as follows. Specifically, both normal accounts and attacked accounts need to provide $N$ audio files to the voice recognition system for training. For a normal account, we randomly split the $N$ audio files into two groups, each with $\frac{N}{2}$ files. Then, we perform the t-SNE analysis on the two groups to identify their similarity. As for the attacked account, we group $\frac{N}{2}$ victim’s audio files together and compare them with $\frac{N}{2}$ attacker’s files using the t-SNE. The comparison results of the two cases are shown in Figure 3 (a) and (b), respectively. As we can see, there is not an obvious difference in the data distribution between the normal account and the attacked account. That means comparing raw audio files may not be an effective way to detect the attackers.
@@ 5 Experimental Studies
In the experiments, we use two datasets: LibriSpeech [75], VoxCeleb [76, 77, 78]. The LibriSpeech and VoxCeleb datasets contain audio files from 2338 and 7245 people, respectively. More statistics can be found in Table I and II. For LibriSpeech, we combined the ”train-clean-100” and ”train-clean-360” partitions for training, and used the ”train-other-500” partition for evaluation. All LibriSpeech partitions were trimmed down to contain 10 audio files per user; Both systems were also trained with the VoxCeleb2 dataset, and evaluated with VoxCeleb1. For Deep Speaker, we again trimmed the two VoxCeleb datasets to contain 10 files per user; however, for the VGG Model, we kept the original VoxCeleb2 dataset for training, and the trimmed VoxCeleb1 dataset for evaluation. 10 Deep Speaker models are trained by varying the weight initialization and the set of users being attacked. We evaluate the scenarios when the percentages of attacked users are 5% or 10%. For the Guardian network, 10 interleaved feature vectors are generated for each account and the value $K$ is set to 11 for the KNN classifier. To choose the optimal $K$ value, we first plot the curve of the error rate and $K$ with $K$ varying from 1 to 30. Then, according to the graph, we select the $K$ value with the projected minimum error rate.
@@
All the experiments were conducted on a computer with Intel i9-10900X CPU@3.7GHz, NVIDIA GeForce RTX 3090 GPU, and 64GBs of memory. Our Guardian model takes about 25 minutes for training. During evaluation mode, the Deep Speaker model takes 0.954 seconds to register a user, and the Guardian model takes 0.243 seconds to validate a user. In what follows, we focus on evaluating its effectiveness in terms of prediction accuracy, recall, and false positive rate (FPR) as defined in the following equations.
@@
$FPR=\frac{Incorrectly~{}~{}Predicted~{}~{}Normal~{}~{}~{}Accounts}{Total~{}~{}Number~{}~{}of~{}~{}Normal~{}~{}Accounts}$
5.1 Comparison with Existing Defense Mechanisms
At the beginning, we aim to compare the performance of our Guardian network with defense mechanisms that employ conventional machine learning algorithms and the latest data poisoning defense mechanism that used fully connected layers (denoted as ”FC” in the figures) [16]. Table V shows the attack detection accuracy, recall and false positive rate when there are 5% and 10% of poisoned user accounts. In both cases, we observe that our proposed Guardian network has achieved around 95% detection accuracy and recall, whereas other approaches have less than 60% accuracy and lower recall. The recall of the Guardian is also much higher than the other three approaches. These can be attributed to the series of techniques adopted by the Guardian network. The results clearly demonstrate the significant benefits of bias reduction, input augmentation, convolutional layers and ensemble learning.
TABLE V: Datasets with 5% and 10% Attacker Accounts
Poisoning Ratio Method Accuracy Recall FPR
5% Attacker Accounts SVM 0.543 0.754 0.468
KNN 0.520 0.583 0.483
FC 0.639 0.804 0.370
Guardian 0.944 0.952 0.057
10% Attacker Accounts SVM 0.488 0.452 0.508
KNN 0.520 0.528 0.481
FC 0.529 0.588 0.478
Guardian 0.939 0.910 0.058
@@ 5.2 Effect of Bias Reduction
In this round of experiments, we are interested in examining the effect of our bias reduction technique alone. Table VI shows the detection accuracy and recall with and without the bias reduction technique. The percentage of poisoned accounts is set to 5%. The version without bias collects training data from only one Deep Speaker model and has the ratio of the attacked accounts to the normal accounts as 1:20. The version with bias reduction used 4 Deep Speaker models and increased the ratio of the attacked accounts to the normal accounts to 1: 5.
TABLE VI: Effect of Bias Reduction
Bias Reduction Accuracy Recall FPR
No 0.960 0.266 0.006
Yes 0.944 0.952 0.057
@@ 5.3 Effect of Input Augmentation
We also evaluate the effectiveness of our proposed input augmentation, which interleaves a pair of feature vectors from the same account. Specifically, we modify the Guardian network to take the 512-dimensional feature vectors generated by the Deep Speaker model as input. We still keep the bias reduction and ensemble learning techniques for the 512-dimensional Guardian network so as to single out the effect of the input augmentation. We compare this modified Guardian network with the Guardian network that has the input augmentation and takes 1024-dimensional feature vectors.
Table VII reports the performance of these two networks. We can observe that a significant performance improvement has been achieved by the input augmentation technique for both the accuracy and recall. Specifically, the accuracy has been increased from 66% to 95%, and the recall has been increased from 35% to 95%. This is because when 512-dimensional feature vectors are used, the convolutional layers only study the spatial relationship in a single vector. The interleaved 1024-dimensional feature vectors give the convolutional layers an opportunity to compare the features from the attacker and victim dimension by dimension, and hence lead to better classification capabilities.
TABLE VII: Effect of Input Augmentation
Input Type Accuracy Recall FPR
512-D (Original) 0.663 0.350 0.321
1024-D (Augmented) 0.944 0.952 0.057
@@ 5.4 Effect of Ensemble Learning
In Table VIII, we show the performance of two versions of the Guardian network with and without ensemble learning. The one without ensemble learning conducts only one prediction by taking a single 1024-dimensional feature vector from a randomly selected pair of 512-dimensional feature vectors from a new user account. The one with the ensemble learning conducts 10 predictions from 10 randomly selected pairs of the 512-dimensional feature vectors from the same user account.
@@
TABLE VIII: Effect of Ensemble Learning
Ensemble Learning Accuracy Recall FPR
No 0.866 0.458 0.113
Yes 0.944 0.952 0.057
@@ 5.5 Effect of Attackers’ Voices
In the previous experiments, both the victims and attackers are chosen randomly, which means the voices of the victim and the attacker could be relatively similar (e.g., of the same gender) or very different. The attack has been successful in both cases. We are interested in finding out if an attacker who has a similar voice as the victim would impose more challenges on Guardian. In this experiment, we use the Deep Speaker and Guardian models trained for Section 5.3. Then, we test Guardian against two kinds of attackers. Specifically, we select attackers of the same gender as the victims to simulate the similar voice scenario. Table IX compares the detection accuracy, recall and false positive rate (FPR) of Guardian under different attack scenarios. We can observe that the Guardian’s performance is not affected by the gender between attacker and victim’s voices.
TABLE IX: Effect of Attacker’s Gender
Attacker’s Voices Accuracy Recall FPR
Mixed 0.944 0.952 0.057
Same Gender 0.942 0.967 0.059
Different Gender 0.942 0.976 0.060
Next, we examine more challenging scenarios when the attackers’ voices are extremely similar to the victims based on two commonly used similarity metrics: Consine similarity and clustering metrics. Such attackers are chosen as follows:
  1. 1.
    Cosine Similarity: Each user account has 10 unique voice files; through the Deep Speaker model, we can obtain 10 corresponding feature vectors. We then calculate the cosine similarities between each pair of feature vectors, resulting in a total of 45 unique comparisons. Finally, by calculating the mean value of these 45 pairs, we can obtain a single cosine similarity score that represents the overall similarity of the 10 voice files. Based on the results, a threshold for cosine similarity scores is established to allow majority (95%) of normal accounts to be classified as acceptable. If the cosine similarity score between an attacker and the victim is above this threshold, such attacker is considered to have extremely similar voices to the victim and added to the test dataset for our Guardian model to check.
  2. 2.
    Clustering Metrics (LDA): We employ the clustering metrics and within-class scatter matrices concept to define “similar voices”. More specifically, we use Linear Discriminant Analysis (LDA) to compress the original 512-dimensional feature vector into a smaller 150-dimensional feature vector, the same dimension as the x-vector model [80]. We then calculate the concentration of within-class feature vectors using the following formula:
    $m_{i}=\frac{1}{N}\sum_{x\subset User_{i}}^{N}x_{k}$ (3)
    $S_{i}=\sum_{x\subset User_{i}}^{N}(x-m_{i})(x-m_{i})^{T}$ (4)
    $ConcentrationScore_{User_{i}}=Trace(S_{i})$ (5)
    A threshold for concentration scores is established to allow the majority (95%) of normal accounts to be classified as acceptable. If an attacker ’s concentration score is below this threshold, the attacker is added to the test dataset for the Guardian model to check.
Table X shows the percentage of such extremely similar attackers that can be detected by our Guardian model. Observe that among all the attackers which can evade the detection using the cosine similarity threshold, 27% can still be captured by Guardian; among all the attackers whose concentration scores to the victims are below the clustering metric threshold, around 60% can still be captured by Guardian. This results indicate the superiority of our Guardian model to approaches that applies simple similarity metrics. It also points out future research directions on further enhancing Guardian model on detecting such extremely similar voices.
TABLE X: Detecting Attackers with Extremely Similar Voices to the Victims
Similarity Metrics Percentage of Attackers Detected
Cosine Similarity 27%
Clustering Metrics 60%
5.6 Effect of Datasets
In this round of experiments, we are interested in learning how audio files with noises would affect the performance of Guardian. It is worth noting that for most of voice authentication, audio files are expected to contain less noises. This experiment just aims to test the limit of our approach.
Table XI shows the performance Guardian when detecting poisoned Deep Speaker model that is trained on LibriSpeech and VoxCeleb, respectively. The VoxCeleb dataset contains lots of background noises compared to LibriSpeech, and the Guardian model’s detection accuracy in VoxCeleb dataset is lower than that in the LibriSpeech dataset. This is because the voice recognition accuracy of the Deep Speaker with respect to VoxCeleb dataset is already lower than that with respect to LibriSpeech as shown earlier in Section 3.3. As a result, this low voice recognition accuracy (88%) of the Deep Speaker affects the detection accuracy of the Guardian model. To sum up, the Guardian model will perform well when the corresponding voice recognition model performs well.
TABLE XI: Model Transferability
Dataset Accuracy Recall FPR
Deep Speaker with LibriSpeech 0.944 0.952 0.057
Deep Speaker with VoxCeleb 0.863 0.831 0.135
5.7 Detecting Pure Man-in-the-Middle Attack
Finally, we examine the capability of our Guardian model in terms of detecting pure man-in-the-middle (MiTM) attacks. As shown in Table 4 in Section 3.3, pure MiTM attacks have success attack rate around 88%. Table XII shows the performance of our Guardian model for the poisoned and clean authentication models. We observed that the Guardian model could achieve even better performance when the voice authentication model has not been poisoned. This also indicates that it would be easier for attackers to impersonate the victims if the attackers can poison the model during the training phase. On the other hand, poisoned data leads to bad feature vectors which imposes greater challenges to the Guardian model due to the less clear differences between the two feature vectors as compared to those generated by a clean model. Applying the Guardian to the clean model eliminates the adverse effects of data poisoning attacks, which explains the better detection accuracy in this experiment.
TABLE XII: Success Rate of Detecting Pure Man-in-the-Middle Attacks
Dataset Accuracy Recall FPR
Poisoned Deep Speaker (LibriSpeech) 0.944 0.952 0.057
Clean Deep Speaker (LibriSpeech) 0.970 0.981 0.031
Poisoned VGG (LibriSpeech) 0.953 0.939 0.046
Clean VGG (LibriSpeech) 0.963 0.939 0.036
@@ 6 Security Analysis
We now analyze potential attacks to our Guardian model. The first scenario that may seem to challenge the effectiveness of the Guardian model is if the attacker has a similar voice to the victim. The similarity here simply refers to how the voice sounds to the human ears. Fortunately, the voice recognition models have much better voice identification capabilities than humans. They already have very high recognition accuracy among a large number of users whereby users of similar voices inevitably exist. In our experiments (Section 5.5), we also simulated such a scenario by selecting same-gender attacker and victims since their voices would be more similar than those from different genders. The experimental results show that our Guardian model performs similarly well in both cases. In the future, to further investigate the performance of the Guardian, we plan to train our Guardian model with “similar voices” instead of randomly selected attacked users.
TABLE XIII: Voice Recognition Accuracy Under Fake Voice Attacks
@@
Another scenario is when the attacker knows the existence and architecture of our Guardian model and attempts to take advantage of that. The attackers may try to use some existing Text-to-Speech (TTS) techniques to generate voice files that mimic the victim’s voices so as to fool the voice authentication model and the Guardian model. However, even if the fake voices are successfully generated to fool the voice authentication model and the Guardian model, they will not be able to escape from the fake voice detector which is good at distinguishing fake voices from authentic human voices. We have conducted the following experiments to validate our conjecture. First, we used a well-known repository called Real-Time Voice Cloning [81] to generate fake voices. This repository is an implementation of SV2TTS [82]. There are three different neural network structures in this system, and we use the pre-trained models they provided during our experiments. We chose 1166 users in LibriSpeech [75] dataset. Keeping the same settings as that in the previous experiments, each user has 10 original voice files, and 5% of user accounts (i.e., victims) contain fake voice files. Each victim account has 5 original voice files and 5 fake voice files generated by the Real-Time Voice Cloning system. Table XIII shows the accuracy of voice recognition before and after the data poisoning attack. Before the attack, the overall voice recognition accuracy is around 0.991. After some user accounts being injected with fake voices, the recognition accuracy of both normal accounts and victim accounts still stay at a very high level as shown in the table. That means the fake voices have successfully fooled the voice recognition model. Next, we check how the Guardian model reacts to the fake voices. Table XIV shows the detection accuracy and recall. From the low recall rate, we can see that the Guardian model misses the majority of fake voices.
TABLE XIV: Guardian with 5% Fake Voice Users
Poisoning Ratio Accuracy Recall FPR
5% Fake Voice 0.953 0.136 0.005
@@
The results from Table XIII and XIV indicate that both the voice recognition model and the Guardian model are unaware of the fake voice attack. However, there is an easy way for the service providers to filter out such fake voices by employing the existing fake voice detectors such as Deep Sonar [79]. This detector is based on monitoring neuron behaviors of the voice recognition system to discern synthesized fake voices. In order to observe the performance of Deep Sonar in our system, we trained a brand new Deep Sonar network for the Deep Speaker using the LibriSpeech dataset. Table XV shows the high detection accuracy and recall achieved by Deep Sonar with respect to the fake voice attack. This means Deep Sonar is a very effective tool to detect fake voices and is powerful enough to defend against the fake voice attack.
@@
Note that fake voice detectors are a good complement to the overall voice authentication system, but cannot replace the function of the Guardian model since fake voice detectors are only versed at identifying manipulated voices. The second row in Table XV shows nearly zero recall rate with respect to poisoning attacks using human voices. In other words, Deep Sonar is not able to filter out authentic human voices when the attacker directly injected their real voices in the victim’s account like the targeted data poisoning attack in our case.
TABLE XV: DeepSonar under Different Types of Attacks
Attack Name Accuracy Recall FPR
5% Fake Voice Attack 0.941 0.939 0.059
5% Poisoning Attack 0.891 0.05 0.065
@@
We also consider the scenario when attackers attempt to apply the idea of Generative Adversarial Networks (GAN) [83] to produce voice files that can fool the voice authentication model, the fake voice detector, and the Guardian model. However, it would be extremely challenging to implement such an attack. Recall that both the fake voice detector and the Guardian model hide behind the voice recognition model. As the attacker does not have the specific parameters of any of these three models, they may train their own system with the same structures. Since all the three models are deep neural networks, the back propagation process is very complicated. In our trial with known system parameters, we are still not able to make such training converge. Even if the attackers managed to complete the local training, it is unclear if the locally generated fake voices will be sufficiently effective to fool the real models used by their targeted service providers.
Finally, we consider limitations around the specific attack and threat model. While our threat model includes a data poisoning attack and a MiTM attack, since this area of research is still relatively new, other sophisticated attacks may emerge as well. For example, [27] analyzes data poisoning attacks in situations where the model is periodically re-trained whereas in our work, the voice authentication system is assumed not to be retrained any more. Additionally, our Guardian model is designed to protect CNN-based voice recognition models. It is worth studying whether Guardian may also be effective for other architectures like transformers and wav2vec [84].